# Tally: Non-Intrusive Performance Isolation for Concurrent Deep Learning Workloads

### v1

@@ 4.2. Priority-aware Scheduling
We now demonstrate how Tally’s priority-aware scheduler leverages the aforementioned slicing and preemption primitives and the transparent profiler to enforce priority among the scheduling of concurrent workloads. The algorithm of the scheduler, as outlined in Figure 3, follows an opportunistic strategy: it prioritizes the immediate dispatch of high-priority kernels upon their arrival (lines $14$ to $20$) and schedules low-priority kernels to execute when the high-priority job becomes inactive (lines $21$ to $33$). Due to the reactive nature of this algorithm, incoming high-priority kernels may need to wait for the completion of currently executing low-priority kernels, potentially introducing delays to the execution of high-priority tasks. To mitigate such queuing delays, each best-effort kernel is either launched as multiple sub-kernels one at a time, or in its preemptive version, such that they can be preempted in the event of a high-priority kernel arriving during its execution.
@@ 4.3. Non-intrusive Virtualization Layer
Finally, we present details on the implementation of Tally’s virtualization layer. On the application-side, Tally utilizes the LD_PRELOAD mechanism in Linux to intercept device API calls, redirecting them to a surrogate function that mirrors the API’s original signature. These intercepted API calls are sent to the server, which handles the actual device execution. The challenge in this process is enabling the server, which operates in a separate address space, to execute client kernel functions. The solution leverages a critical insight: GPU applications register device code with the GPU driver at the start of execution. By intercepting this process, the server gains access to the device code, including kernel binaries and device code, which we use at the server-side to apply kernel transformations and recompile to executable GPU kernels.
@@ 5.1. Methodology
Workloads. For Tally evaluation, we prepare a benchmark suite comprising six training and six inference workloads as listed in Table 2, covering a diverse range of widely used DL model architectures, such as CNNs (e.g., ResNet50 (kaiming2016, )), transformer-based LLMs (e.g., Llama-2 (touvron2023llama, )), and diffusion models (e.g., Stable Diffusion (rombach2022high, )). Training workloads are implemented using PyTorch (paszke2019pytorch, ), a broadly adopted DL framework in the research community. For inference workloads, we use high-performance engines such as ONNX Runtime (onnxruntime, ), TorchInductor (torchinductor2024, ), and Hidet (hidet2023, ). Inspired by prior DL serving benchmarks (li2023-alpaserver, ; reef2022, ; bhattacharjee2019barista, ), we simulate inference workloads using the Microsoft Azure Function Trace 2021 (MAF2) (maf2, ), a dataset of Azure serverless function invocations that has been repurposed for DL serving workloads. We extract the trace of the most frequently invoked function to adapt it to match specific traffic loads based on the query latency of the models. We define load as the percentage of time an inference service is actively serving requests.
Metrics. We use the $99^{th}$ percentile latency as the primary metric for assessing the performance of inference tasks. For both training and inference tasks, we measure their throughput, defined as the number of samples processed per unit of time. To evaluate the overall performance of concurrent task execution, we utilize system throughput, calculated by aggregating the normalized throughput of individual workloads.
@@
Baselines. We compare the performance of Tally against four state-of-the-art non-intrusive GPU sharing solutions: (i) Time-Slicing, (ii) MPS, (iii) MPS-Priority and (iv) TGS. All these solutions perform scheduling at the kernel-level granularity. More specifically, Time-Slicing and MPS are the proprietary concurrency solutions provided by NVIDIA GPUs. MPS-Priority refers to MPS with the client priority control feature enabled, which allows specifying priority levels among concurrent workloads. Lastly, TGS is a GPU sharing system specifically designed for DL workloads, which leverages adaptive rate control mechanisms to achieve performance isolation among concurrent workloads.
Experimental setup. All experiments are conducted on an AWS EC2 p4d.24xlarge instance, a production-grade GPU server widely deployed in DL clusters. This machine has $96$-core Intel Xeon Platinum 8275CL CPU, $1152$ GB of DRAM, and $8$ NVIDIA A100 GPUs, each with $40$ GB of memory. The server operates on Ubuntu 20.04 with CUDA 12.2, PyTorch 2.2.0, ONNX Runtime GPU 1.17.0, and Hidet 0.3.0.
5.2. End-to-end Results
We first analyze Tally’s end-to-end performance in efficiently sharing GPU resources among multiple workloads while guaranteeing performance isolation for high-priority inference tasks. To this end, we measure the $99^{th}$-percentile latencies of the inference tasks and the overall system throughput achieved by Tally when executed concurrently with a best-effort training workload. First, we compare the latency metrics against the state-of-the-art non-intrusive GPU sharing systems, namely Time-Slicing, MPS, MPS-Priority, and TGS, as well as against the ideal scenario of running the inference task in isolation. Figure 4 presents the latency and throughput measurements for all inference-training combinations from the benchmark suite, with inference tasks simulated using the MAF2 trace with a fixed $50\%$ load. The results demonstrate Time-Slicing, MPS, MPS-Priority, and TGS exhibit significantly high overhead compared to the ideal latency, with an average of $252.3\%$, $345.0\%$, $195.5\%$, and $188.9\%$, respectively. The suboptimal inference latencies exhibited by Time-Slicing and MPS can be attributed to their priority-agnostic design, which focuses solely on maximizing system utilization. MPS-Priority and TGS, on the other hand, exhibit improved latency by prioritizing kernel execution of high-priority tasks over best-effort tasks. However, their coarse-grained kernel-level scheduling strategy fails to effectively mitigate co-execution interference. In stark contrast, Tally consistently maintains tail latencies nearly identical to the ideal latencies of the high-priority workloads, with an average percentage increase of mere $7.2\%$. This stronger performance isolation capability can be attributed to the block-level scheduling strategy that we follow.
Furthermore, we observe that MPS, MPS-Priority, and TGS exhibit substantial variation in performance degradation for high-priority inference tasks across different workload combinations. For instance, TGS experiences performance degradation ranging from $15.6\%$ to $751.7\%$. This variation in latency reveals the sensitivity of these systems to workload characteristics, such as the differences in kernel duration across various tasks. As these systems schedules at the kernel-level, their performance deteriorates more severely for workloads with longer kernels. In contrast, Tally’s block-level scheduling demonstrates robust performance against variations in kernel duration, effectively limiting the performance degradation of inference tasks to under $10\%$ across $70\%$ of the workload combinations (worst-case $23\%$ on ResNet50).
@@ 5.3. Traffic Load Sensitivity Analysis
To further analyze the adaptive nature of Tally to varying traffic load, in Figure 5v, we plot the real-time traffic trace to a BERT inference task (a condensed version of the original MAF2 trace), and measure the real-time tail latency and throughput when co-located with another BERT training task. Figure 5v illustrates (1) the real-time traffic to a BERT inference task, (2) $99^{th}$ percentile latency of the inference server achieved by different GPU sharing systems and (3) the throughput of the co-located best-effort BERT training achieved by Tally compared to the original throughput of the training task. As shown in Figure 5b (2), the tail latency under Tally, represented by the purple line, closely aligns with the ideal latency depicted by the blue line, with negligible deviation throughout the simulation period. Whereas alternative approaches such as Time-Slicing, MPS, MPS-Priority, and TGS result in substantial slowdowns in latency as depicted. Additionally, Figure 5b (3) shows Tally’s ability to opportunistically adjust the throughput of the best-effort task in correspondence with the fluctuating traffic load to the high-priority inference task, as discussed in Section 4. As a result, Tally preserves over $68\%$ of the training task’s original throughput simultaneously running the high-priority inference workload on the same GPU without affecting its latency requirements.
@@ 5.5. Ablation study
Performance decomposition. We analyze the importance of Tally’s priority-aware scheduling algorithm and block-level transformation for performance isolation. Figure 6(b) compares the $99^{th}$-percentile latency of a high-priority BERT model when co-located with training workloads from the benchmark suite. Ideal denotes the original latency of the task executed in isolation. No-scheduling refers to the indiscriminate dispatching of kernels from both workloads upon arrival. Scheduling w/o transformation employs Tally’s priority-aware scheduling without kernel transformations, while Scheduling with transformation represents the full implementation of Tally, utilizing slicing and preemption mechanisms for block-level scheduling. The results reveal that No-scheduling incurs substantial latency overhead across all six job combinations, with the most severe case exhibiting a 30×30$\times$30 × slowdown when co-located with Whisper. This significant performance degradation is attributable to the absence of priority enforcement, resulting in considerable co-execution interference. The integration of Tally’s priority-aware kernel-level scheduling strategy somewhat mitigates this overhead. For ResNet50 and GPT2, this approach yields latencies close to ideal performance, with $8.0\%$ and $9.8\%$ slowdown respectively. However, for workloads such as Whisper and BERT, the latency slowdown can still reach up to 10×10$\times$10 ×. This difference in slowdown is attributed to the variation in kernel duration distribution. For instance, $99.3\%$ of kernels in ResNet50 complete in less than $0.1$ ms, whereas $5.6\%$ of kernels in Whisper exceed the duration of an entire BERT inference ($3.93$ ms). These prolonged kernels significantly delay the execution of high-priority kernels. Notably, with kernel transformation enabled, i.e., with full Tally, the latency overhead becomes negligible across all six training workloads, resulting in an average of $4.0\%$ slowdown ($6.2\%$ in the worst case). This shows the necessity of block-level scheduling in maintaining robust performance isolation guarantees.
@@
Turnaround latency analysis. We examine the impact of the turnaround latency threshold (delineated in Section 4) by evaluating the tail latency of the BERT inference workload and the normalized throughput of the co-located training job on six models and across six distinct threshold values, ranging from $0.01$ ms to $10$ ms. The results, depicted in Figure 6 (c-i) and (c-ii), show that higher turnaround latency thresholds generally correspond to increased tail latency for the inference workload with only slight increase in the throughput. Based on these empirical results, we find that a turnaround latency threshold of $0.0316$ ms provides the best balance between latency and throughput, which is the default setting we use throughout experiments.

### v3

@@ 4.2. Priority-aware Scheduling
We now demonstrate how Tally’s priority-aware scheduler leverages the aforementioned slicing and preemption primitives alongside the transparent profiler to enforce priority during the scheduling of concurrent workloads. The algorithm of the scheduler, as outlined in Figure 4, follows an opportunistic strategy: it prioritizes the immediate dispatch of high-priority kernels upon their arrival (lines $14$ to $20$) and schedules low-priority kernels to execute only when the high-priority job becomes inactive (lines $21$ to $33$). Due to the reactive nature of this algorithm, incoming high-priority kernels may need to wait for the completion of currently executing low-priority kernels, potentially introducing delays to the execution of high-priority tasks. To mitigate such queuing delays, each best-effort kernel is either launched as multiple sub-kernels one at a time, or in its preemptive form, such that it can be preempted in the event of a high-priority kernel arriving during its execution.
@@
To strike a balance between ensuring robust performance isolation for high-priority tasks and maintaining acceptable throughput for low-priority tasks, Tally’s priority-aware scheduler employs a profile-guided resource provisioning policy. This policy leverages runtime performance metrics to determine the optimal launch configuration for each best-effort kernel through a search process. As depicted in Figure 4, lines $2$ to $10$, for each kernel, the scheduler generates a set of candidate configurations encompassing both slicing and preemption. For preemption, potential configurations are derived by considering all possible multiples of the number of SMs that fit with the thread constraints. Whereas slicing may set the degree of slicing based on having each slice operating on different percentages of the total blocks. These candidate configurations are then assessed by Tally’s transparent profiler.
To evaluate the performance of candidate configurations, Tally employs the turnaround latency metric, as introduced in Section 3. This metric represents the estimated time needed for a best-effort kernel to release occupied resources. Calculating the turnaround latency for a sliced kernel is relatively straightforward, which can be done by simply measuring the completion time of a single kernel slice. In contrast, determining the turnaround latency for preemptive kernels is more challenging due to their use of dynamic worker blocks. To address this challenge, we employ a heuristic approach that approximates the turnaround latency by dividing the total execution time of the kernel by the number of blocks allocated to each worker, as expressed by:
@@
This formula approximates the latency of waiting for each worker to complete the current block of execution. After profiling each configuration, the scheduler selects the one that achieves the optimal performance while complying with a predefined turnaround latency threshold, which serves as a configurable parameter to balance the performance trade-offs between high and low-priority workloads. Empirically, we find that a default value of $0.0316$ ms can effectively render robust performance isolation while producing competitive system throughput (details in Section 5.6).
Notably, Tally’s profiling process accounts for variations in kernel inputs by profiling each unique work configuration (block and grid dimensions) separately for every kernel. For measurement stability, profiling results are averaged across many runs (e.g., $10$). Since profiling operates on a per-kernel basis, once measurements for a kernel are collected, they can be reused throughout the execution, thus introducing negligible overhead to overall execution.
@@ 4.3. Non-intrusive Virtualization Layer
Finally, we present details on the implementation of Tally’s virtualization layer. On the application-side, Tally utilizes the LD_PRELOAD mechanism in Linux to intercept device API calls, redirecting them to a surrogate function that mirrors the API’s original signature. These intercepted API calls are sent to the server, which handles the actual device execution. The challenge in this process is enabling the server, which operates in a separate address space, to execute client kernel functions. The solution leverages a critical insight: GPU applications register device code with the GPU driver at the start of execution. By intercepting this process, the server gains access to the device code, including kernel binaries and PTX code, which we use at the server-side to apply kernel transformations and recompile to executable GPU kernels.
@@ 5.1. Methodology
Workloads. To evaluate Tally, we prepare a benchmark suite consisting of six training and six inference workloads as listed in Table 2, covering a diverse range of widely used DL model architectures, including CNNs (e.g., ResNet50 (kaiming2016, )), transformer-based LLMs (e.g., Llama-2 (touvron2023llama, )), and diffusion models (e.g., Stable Diffusion (rombach2022high, )). Training workloads are implemented using PyTorch (paszke2019pytorch, ), a broadly adopted DL framework in the research community. For inference workloads, we use high-performance inference engines such as ONNX Runtime (onnxruntime, ), TorchInductor (torchinductor2024, ), and Hidet (hidet2023, ). Some kernels in the workloads are sourced from proprietary libraries such as cuBLAS (cuBLAS, ), which restrict device code interception. To enable block-level transformations, Tally automatically replaces these kernels at runtime with CUTLASS (cutlass, ) alternatives that offer similar performance. Inspired by prior DL serving benchmarks (li2023-alpaserver, ; reef2022, ; bhattacharjee2019barista, ), we simulate inference workloads using the Microsoft Azure Function Trace 2021 (MAF2) (maf2, ), a dataset of Azure serverless function invocations that has been extensively repurposed for DL serving research. We extract the trace of the most frequently invoked function and adapt it to match specific traffic loads based on the inference latency of the models. We define load as the percentage of time an inference service is actively serving requests.
Metrics. We use the $99^{th}$ percentile latency as the primary metric for assessing the performance of inference tasks. For both training and inference tasks, we measure their throughput, defined as the number of samples processed per unit of time. To evaluate the overall performance of concurrent task execution, we calculate system throughput by summing the normalized throughput of all concurrent workloads.
Baselines. We compare the performance of Tally against four state-of-the-art non-intrusive GPU sharing solutions: (i) Time-Slicing, (ii) MPS, (iii) MPS-Priority, and (iv) TGS. All these systems perform scheduling at the kernel-level granularity. Specifically, Time-Slicing and MPS are the proprietary concurrency mechanisms provided by NVIDIA GPUs. MPS-Priority refers to MPS with the client priority control feature enabled, which allows assigning priority values to workloads for prioritized resource allocation. TGS is a GPU sharing system specifically designed for DL workloads that leverages adaptive rate control mechanisms to achieve performance isolation across concurrent workloads.
Experimental setup. All experiments are conducted on an AWS EC2 p4d.24xlarge instance, a production-grade GPU server widely deployed in DL clusters. It is equipped with a $96$-core Intel Xeon Platinum 8275CL CPU, $1152$ GB of DRAM, and $8$ NVIDIA A100 SXM GPUs, each with $40$ GB of memory. The server runs on Ubuntu 20.04, and is configured with CUDA 12.2, PyTorch 2.2.0, ONNX Runtime GPU 1.17.0, and Hidet 0.3.0.
@@
5.2. End-to-end Results
We first analyze Tally’s end-to-end performance in efficiently sharing GPU resources among multiple workloads while guaranteeing performance isolation for high-priority inference tasks. To this end, we measure the $99^{th}$-percentile latencies of the inference tasks and the overall system throughput achieved by Tally when executed concurrently with a best-effort training workload. First, we compare the latency metrics against the state-of-the-art non-intrusive GPU sharing systems, namely Time-Slicing, MPS, MPS-Priority, and TGS, as well as against the ideal scenario of running the inference task in isolation. Figure 5 presents the latency and throughput measurements for all inference-training combinations from the benchmark suite, with inference tasks simulated using the MAF2 trace with a fixed $50\%$ load. The results demonstrate that Time-Slicing, MPS, MPS-Priority, and TGS exhibit significantly higher overhead compared to the ideal latency, with an average increase of $252.3\%$, $345.0\%$, $195.5\%$, and $188.9\%$, respectively. The suboptimal inference latencies exhibited by Time-Slicing and MPS can be attributed to their priority-agnostic design, which focuses solely on maximizing system utilization. MPS-Priority and TGS, on the other hand, exhibit improved latencies by prioritizing kernel execution of high-priority tasks over best-effort tasks. However, their coarse-grained kernel-level scheduling strategies fail to effectively mitigate co-execution interference. In stark contrast, Tally consistently maintains tail latencies nearly identical to the ideal latencies of the high-priority workloads, with an average percentage increase of mere $7.2\%$. This stronger performance isolation capability can be attributed to the block-level scheduling strategy that Tally employs.
@@ 5.2. End-to-end Results
Furthermore, we observe that MPS, MPS-Priority, and TGS exhibit substantial variation in performance degradation for high-priority inference tasks across different workload combinations. For instance, TGS experiences performance degradation ranging from $15.6\%$ to $751.7\%$. This variation in latency reveals the sensitivity of these systems to workload characteristics, such as the differences in kernel duration across various tasks. As these systems schedule at the kernel level, their performance deteriorates more severely for workloads with longer kernels. In contrast, Tally’s block-level scheduling demonstrates robust performance against variations in kernel duration, effectively limiting the performance degradation of inference tasks to under $10\%$ across $70\%$ of the workload combinations (worst-case $23\%$ on ResNet50).
@@ 5.3. Traffic Load Sensitivity Analysis
To illustrate the adaptive nature of Tally to varying traffic load, in Figure 6b, we plot the real-time traffic trace to a BERT inference task (a condensed version of the original MAF2 trace), and measure the real-time tail latency and throughput when co-located with another BERT training task. Figure 6b illustrates (1) the real-time traffic to a BERT inference task, (2) $99^{th}$ percentile latency of the inference server achieved by different GPU sharing systems and (3) the throughput of the co-located best-effort BERT training achieved by Tally compared to the original throughput of the training task. As shown in Figure 6b (2), the tail latency under Tally, represented by the purple line, closely aligns with the ideal latency depicted by the blue line, with negligible deviation throughout the simulation period. Whereas alternative approaches such as Time-Slicing, MPS, MPS-Priority, and TGS result in substantial slowdowns in latency as depicted. Additionally, Figure 6b (3) shows Tally’s ability to opportunistically adjust the throughput of the best-effort task in correspondence with the fluctuating traffic load to the high-priority inference task, as discussed in Section 4. As a result, Tally preserves over $68\%$ of the training task’s original throughput simultaneously running the high-priority inference workload on the same GPU without affecting its latency requirements.
@@ 5.5. Performance Decomposition
We analyze the importance of Tally’s priority-aware scheduling algorithm and block-level transformations for performance isolation. Figure 7(b) compares the $99^{th}$-percentile latency of a high-priority BERT model when co-located with training workloads from the benchmark suite. Ideal denotes the original latency of the task executed in isolation. No-scheduling refers to the indiscriminate dispatching of kernels from both workloads upon arrival. Scheduling w/o Transformations employs Tally’s priority-aware scheduling without kernel transformations, while Scheduling with Transformations represents the full implementation of Tally, utilizing slicing and preemption mechanisms for block-level scheduling. The results reveal that No-scheduling incurs substantial latency overhead across all $6$ job combinations, with the most severe case exhibiting a 30×30$\times$30 × slowdown when co-located with Whisper. This significant performance degradation is attributable to the absence of priority enforcement, resulting in considerable co-execution interference. The integration of Tally’s priority-aware kernel-level scheduling strategy somewhat mitigates this overhead. For ResNet50 and GPT2, this approach yields latencies close to ideal performance, with $8.0\%$ and $9.8\%$ slowdown respectively. However, for workloads such as Whisper and BERT, the latency slowdown can still reach up to 10×10$\times$10 ×. This difference in slowdown is attributed to the variation in kernel duration distribution. For instance, $99.3\%$ of kernels in ResNet50 complete in less than $0.1$ ms, whereas $5.6\%$ of kernels in Whisper exceed the duration of an entire BERT inference ($3.93$ ms). These prolonged kernels significantly delay the execution of high-priority kernels. Notably, with kernel transformation enabled, i.e., with full Tally, the latency overhead becomes negligible across all six training workloads, resulting in an average of $4.0\%$ slowdown ($6.2\%$ in the worst case). This shows the necessity of block-level scheduling in maintaining robust performance isolation guarantees.
5.6. Turnaround Latency Analysis
@@
We examine the impact of the turnaround latency threshold (delineated in Section 4) by evaluating the tail latency of the BERT inference workload and the normalized throughput of the co-located training job on six models across six distinct threshold values, ranging from $0.01$ ms to $10$ ms. The results, depicted in Figure 7 (c-i) and (c-ii), show that higher turnaround latency thresholds generally correspond to increased tail latency for the inference workload with only slight increase in the throughput. Based on these empirical results, we find that a turnaround latency threshold of $0.0316$ ms provides the best balance between latency and throughput, which is the default setting we use throughout experiments.
5.7. Overhead Analysis
Virtualization. We evaluate Tally’s virtualization overhead by measuring the throughput of each workload in the benchmark suite and comparing it to their baseline performance. Results indicate that virtualization incurs an average overhead of only $1\%$, proving minimal performance impact.
Kernel transformation. We evaluate the overhead of Tally’s kernel transformation by profiling $10K$ kernels from best-effort workloads (including same kernel with varying work dimensions). Our analysis reveals an average overhead of $25\%$. Notably, this overhead is exclusive to best-effort kernels. The performance of high-priority jobs remain unaffected.
Profiling. Tally profiles each new kernel from best-effort tasks to determine its optimal launch configuration. Once collected, these measurements are reused throughout the execution. Thus, Tally’s profiling stage for a job generally completes within minutes. This overhead is negligible given that training workloads often run for hours to days.
6. DISCUSSION
Comparison with MIG. NVIDIA MIG (nvidia-mig, ) enables hardware-level GPU partitioning into isolated instances, providing robust performance isolation for concurrent workloads. However, MIG’s coarse-grained partitioning mechanism inherently limits the maximum resource allocatable to individual workloads, which can result in potential performance degradation of high-priority tasks. This limitation stems from MIG’s requirement for configuring resource partitions prior to task execution; any alteration to these partitions necessitates the termination of all running processes, rendering it impractical in production environments where job interruption is typically prohibited. Even when reconfiguration is feasible, this process can incur significant overhead: resetting partitions takes several seconds via command-line instructions, and applications require additional time from seconds to minutes to resume execution. As a result, MIG struggles to promptly adapt resource allocations to meet the dynamic demands of DL workloads, potentially leading to compromised performance of high-priority workloads. Additionally, MIG’s static allocation scheme can result in suboptimal resource utilization, as idle cycles within one partition cannot be dynamically reallocated to others. These limitations prevent MIG from simultaneously achieving high utilization while maintaining the performance of high-priority tasks. In contrast, Tally effectively addresses these challenges by dynamically allocating resources based on real-time resource demands, allowing high-priority tasks to fully utilize the GPU when active while reallocating idle cycles to best-effort tasks, thereby ensuring both high utilization and effective performance isolation.
Generalization to CUDA extensions. While the standard GPU programming model enforces the independence of thread blocks within a kernel (as explained in Section 2), recent CUDA extensions, such as Thread Block Clusters (thread-block-clusters, ) and Cooperative Groups (cooperative-groups, ), introduce support for inter-block dependencies under certain constraints. We illustrate how Tally can reliably detect and handle these cases. (1) Thread Block Clusters organizes thread blocks into clusters, enabling intra-cluster communication via hardware support. These kernels can be identified by special flags passed at kernel launch. In such cases, Tally applies slicing and preemption at the cluster level, rather than at the thread block level. This ensures execution correctness while enabling more fine-grained scheduling. (2) Cooperative Groups introduces new thread grouping mechanisms that extend the capabilities of the standard CUDA programming model. Among them is inter-block synchronization, which enables coordinated execution across thread blocks. Kernels using this feature are launched through the cudaLaunchCooperativeKernel API, making them easily identifiable. Notably, inter-block synchronization requires that all thread blocks within a kernel must be simultaneously scheduled on the GPU. For these kernels, Tally refrains from applying block-level scheduling to ensure compliance with the concurrent execution requirements. In our evaluation, none of the workloads employ inter-block cooperative groups, suggesting that the absence of block-level scheduling for these kernels has a negligible impact on Tally’s overall performance isolation guarantees.